Let f(:,y) = xy - 1

a. Find \bigtriangledown f at (2,1)

b. Find the directional derivative of f at (2,1) in the direction of -i+3j
Answer #1

@ Let f(x) = x² 42_x # t = 2x 42_1 fyn gydy At 12,1) = (2,1)= 4-123 fy (a,0) = 8 2 9 /.. 3i+g; Ah-we . direction = = -itzi Sa